OpenStudy (abbycross167):

Can someone please show me the steps to this question? Please Solve the system of equations: x = 3y – 6 2x – 4y = 8

OpenStudy (abbycross167):

@retirEEd can you try to help me please?

rishavraj (rishavraj):

just plug the value of x in 2nd equation :))

rishavraj (rishavraj):

@abbycross167 there????

OpenStudy (abbycross167):

I am so sorry!!! :( My WIFI messed up... but I'm here now

OpenStudy (abbycross167):

I'm not sure how to "plug the value of x in 2nd equation" like this? x = 3y – 6 = 3y-6=3y? 2x – 4y = 8

rishavraj (rishavraj):

the first equation says x = 3y - 6 so plugging value of x in 2nd eqn means 2(3y - 6) - 4y = 8

OpenStudy (abbycross167):

oh ok

OpenStudy (abbycross167):

@rishavraj what would I do next?

rishavraj (rishavraj):

the equation u got is \[2(3y - 6) - 4y = 8\] solve the equation to get y

OpenStudy (abbycross167):

so this equation is y=10?

OpenStudy (abbycross167):

@rishavraj

rishavraj (rishavraj):

yup :)) good work ))

rishavraj (rishavraj):

now plug tht value in any equation and get x :)

OpenStudy (abbycross167):

Could you show me how to do that @rishavraj please?

rishavraj (rishavraj):

see u got the value of y i.e 10 now x = 3y - 6 so x = 3(10) - 6

OpenStudy (abbycross167):

Thank you very much!!! You helped me very much! :)